1) Company XYZ has a monopoly on a new product. The product can be manufactured in either of two factories. The costs of prodNeed Help Please!
Answer #1

The monopolist's demand curve is

P = 500 – 4Q

The costs of the two factories are

C1 = 5Q21 and C2 = 10Q22

The goal of the Monopolist is to maximize profit

Π = TR – C1 – C2

TR = Total Revenue

TR = PQ

TR = (500 – 4Q)Q

TR = 500Q – 4Q2

Marginal revenue (MR) = ΔTR/ΔQ = 500 – 8Q = 500 – 8(Q1 + Q2)

C1 = 5Q21

MC1 = ΔC1/ΔQ1 = 10Q1

C2 = 10Q22

MC2 = ΔC2/ΔQ2 = 20Q2

Equating each MC to the common MR

500 – 8Q1 – 8Q2 = 10Q1

500 – 8Q1 – 8Q2 = 20Q2

Solving for Q1 and Q2 we find

Q1 = 22.72

Q2 = 11.36

So, the total Q is 34.08 units. This total output will be sold at price P defined by

P = 500 – 4Q

P = 500 – (4 × 34.08)

P = 363.68

The Monopolit’s Profit is

Π = TR – C1 – C2

Π = (363.68 × 34.08) – 5(22.72)2 – 10(11.36)2

Π = 12,394.21 – 2,580.99 – 1,290.50

Π = 12,394.21 – 3,871.49

Π = 8,522.73

Part b) Now it is given that the cost of labor in Factory 1 has increased while the cost of labor in Factory 2 has remained the same. Under this situation, the output at Factory 1 will decline. The quantity produced at Factory 2 will increase.

However, given the fact that while cost has increased in Factory 1 and remained same at Factory 2, so the overall cost of production has in fact increased. As a result, total output will decline. Since, we are dealing with a Monopolist who has market power to determine the price, the rise in cost will lead to increase in the price of the good.